OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (aroub):

-8+3/4-7 =5/-3 y=mx+b (4,-8) -8=5/-3( 4) +b -8,=20/-3+b, then u find b=? y=-8x+? (b)

OpenStudy (anonymous):

could you just give me the answer and ill compare to mine?

OpenStudy (aroub):

ok what did u get?

OpenStudy (anonymous):

4x+ y=6 but i know thats not right

OpenStudy (aroub):

no! :(

OpenStudy (aroub):

ok so the answer is y=-8x-4/3

OpenStudy (anonymous):

how would i reduce that fraction at the end?

OpenStudy (aroub):

it is

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (aroub):

yes sure! :)

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1}=m\] or: \[\frac{y-(-3)}{x-(7)}=\frac{(-8)-(-3)}{(4)-(7)}\] so: \[\frac{y+3}{x-7}=\frac{5}{3}\] Cross multiply terms: \[3(y+3)=5(x-7)\] to get \[3y+9=5x-35\] and rearranging to get \[5x-3y=44\]